Jenny Lupine, the manager for a division that produces a variety of paper products, was considering the divisional manager's request for a sales forecast for a new line of paper napkins. The divisional manager was gathering data so that he could choose between two different production processes

•The first process would have a variable cost of $10 per case produced and fixed costs of$100,000

•The second process would have a variable cost of $6 per case and fixed costs of$200,000.

•The selling price would be $30 per case in either process.Jenny had just received information from the marketing department that projected annual sales of

30,000 cases for the new product. Jenny was reluctant to report this forecast to the divisional

manager. The first process was labor intensive and her friend Ed would be appointed the line

supervisor. If the second process was chosen, Ed and many other workers would be laid

off. After some consideration, Jenny revised the sales projection to 22,000 cases.

Why did Jenny alter the sales forecast?
Answer #1
There is two type of process is involved in production process.
1 is labor intensive and 2 is Machine intensive
There is difference between variable cost and fixed cost of both production process.
Using the formula for determining the BEQ at which total cost of both the production process
would be same irrespective of process used and above that point process having lower variable cost
would result in lower total cost and below that point process having higher variable cost would results in
lower total cost.
BEQ = Difference in fixed cost / Difference in Variable cost
(200000-100000)/(10-6)
100000/4
25000 Units
So we know that if sales volume will be lower than 25000 process 1 Labor intensive would be preferred
Since Janny want to avoid Laid off of Ed her friend and other workers, She has revised the forecasted
Sales figure to 22,000 from 30,000 in that case jobs of many workers will be saved.
Verification:
Cost at 30000 level using process 1 = 30000 x 10 + 100000 = 400000
Cost at 30000 level using process 2 = 30000 x 6 + 200000 = 380000
Cost at 22000 level using process 1 = 22000 x 10 + 100000 = 320000
Cost at 22000 level using process 2 = 22000 x 6 + 200000 = 332000